Factor completely.
2x2 + 40x + 200

Found 2 solutions by faceoff57, jaya_rohith1@yahoo.com:
Answer by faceoff57(82)

2x^2 + 40x + 200
2(x^2 + 20x + 100)
2(x+10)(x+10)

Answer by jaya_rohith1@yahoo.com(33)

Let 2x^2 + 40x +200 =0
x^2 + 20x +100 =0
(x+10)^2 =0
implies x=-10 , -10